OpenStudy (anonymous):

An investor invested a total of $1,600 in two mutual funds. One earned a 7% profit while the other earned a 2% profit. If the investor's total profit was $97, how much was invested in each mutual?

OpenStudy (anonymous):

lets call the amount invested at \(7\%\) "\(x\)" then since the total is \(\$1,600\) the amount invested at \(2\%\) must be \(1,600-x\)

OpenStudy (anonymous):

the profit is \(7\%\) of \(x\) i.e. \(.07\times x\) plus \(2\%\) of \(1600-x\) or \(.02(1600-x)\) for a total of \[.07x+.02(1600-x)\]

OpenStudy (anonymous):

since you know the total is \(\$97\) you can set \[.07x+.02(1600-x)=97\] and solve for \(x\)

OpenStudy (anonymous):

probably easiest to get rid of the annoying decimals by multiplying both sides by \(100\) and solve \[7x+2(1600-x)=9700\]

OpenStudy (anonymous):

Im lost, What would i do next?

OpenStudy (anonymous):

distribute the \(2\)

OpenStudy (anonymous):

So what is the amount invested in the mutual funds that earned 7% $ ?

OpenStudy (anonymous):

@Jamierox4ev3r could you please help me out with this problem?

HanAkoSolo (jamierox4ev3r):

well, @satellite73 had the process right, but you still need to distribute the equation: 7x+2(1600-x)=9700 7x+3200-2x=9700 ----------------- 5x+3200=9700 -3200 -3200 ---------------- 5x=6500 /5 /5 ------ x=1300

OpenStudy (anonymous):

Thank you very much. So what was the amount invested in the mutual fund that earned 7%?

HanAkoSolo (jamierox4ev3r):

thats right :D

HanAkoSolo (jamierox4ev3r):

lol sorry i was doing another problem but yes thats right

HanAkoSolo (jamierox4ev3r):

@flower93